OpenStudy (helpblahblahblah):

Plz help Milo is a car salesman. He earns an annual salary of $23,000 plus 7% commission on all sales. How much does Milo make in a year that he has $958,000 in car sales?

OpenStudy (anonymous):

for milo annual salary 23,000 now he also earn commission on all sale as 7% his total sales 958000 now its 7% = 958000 * 7/100 = 958000 * 0.07 =67060 so total earning = 23000+67060 = $90,060

OpenStudy (helpblahblahblah):

thanks

OpenStudy (anonymous):

ur welcome ...